# Neural Architecture Search via Two Constant Shared Weights Initialisations

### v2

@@ Abstract
In recent years, zero-cost metrics are gaining ground in neural architecture search (NAS). These metrics allow finding the optimal neural network for a given task faster and with a lesser computational load than conventional NAS methods. Equally important is that they also shed some light on the internal workings of neural architectures. This paper presents a zero-cost metric that highly correlates with the train set accuracy across the NAS-Bench-101, NAS-Bench-201 and NAS-Bench-NLP benchmark datasets. We evaluate a neural architecture’s potential based on the outputs’ statistics after two constant shared weights initialisations. For this, we only use an unlabelled mini-batch of data. We observe that the dispersion of the outputs between two initialisations positively correlates with trained accuracy. The correlation further improves when we normalise dispersion by average output magnitude. The resulting metric, epsilon, does not require gradients computation and unbinds the NAS procedure from training hyperparameters, loss metrics and human-labelled data. Our method is easy to integrate within existing NAS algorithms and takes a fraction of a second to evaluate a single network. The code supporting this study can be found on GitHub at https://github.com/egracheva/epsinas.
Keywords: machine learning, neural architecture search, zero-cost NAS
@@ 1 Introduction
The field of neural architecture search (NAS) emerged about a decade ago as an effort to automatise the process of neural geometry optimisation. At the early stages of NAS development, every candidate architecture used to be evaluated through the training process (reinforcement learning (Williams (1992)), evolutionary algorithms (Real et al. (2019)), Bayesian optimisation (Falkner et al. (2018); White et al. (2021))).
One-shot algorithms adopting weight sharing dispense of multiple architectures training, reducing the search time drastically (efficient reinforcement learning (Pham et al. (2018)), random search with parameters sharing (Li and Talwalkar (2020)), differentiable methods). Nevertheless, they require the training of a massive hypernet, which necessitates elaborate hyperparameter tuning. While these methods prove efficient, they do not systematically achieve satisfactory results (Dong and Yang (2019)). One of the best of them, DARTS- (Chu et al. (2020)), shows significant uncertainty compared to evolutionary or reinforcement algorithms.
There are available methods that estimate network performance without training the dataset of interest but relying on an auxiliary predictive machine learning (ML) model built on a dataset of trained architectures (Istrate et al. (2019); Deng et al. (2017)). These methods aim to accelerate the NAS process for image recognition but still rely on training and cannot apply to other ML problems.
@@
1.1 Zero-cost NAS
To alleviate the process of architecture search, many researchers focus on developing methods that allow finding optimal architectures without model training – so-called zero-cost NAS methods. These methods evaluate networks via a trainless proxy. They typically require the equivalent of one or a few training epochs, which is two to three orders of magnitude faster than previous NAS methods.
Weight agnostic neural networks. One of the pioneering works in zero-shot NAS is presented by Gaier and Ha (2019). The authors demonstrate a constructor building up neural architectures based on the mean accuracy over several initialisations with constant shared weights and the number of parameters contained within the model. The resulting model achieves over $90\%$ accuracy on MNIST data (LeCun et al. (2010)) when the weights are fixed to the best-performing constants. While these results are very intriguing, the authors admit that such architectures do not perform particularly well upon training. Moreover, back in $2019$, the benchmark databases of trained architectures, which are now routinely used to compare NAS metrics with each other, were yet to be released, which disables the comparison of this zero-shot method against the most recent ones.
@@ 1.1 Zero-cost NAS
ReLU activation pattern. In $2020$, Mellor et al. (2020) present the naswot metric, which exploits the rectified linear unit (ReLU, Agarap (2018)) activation function’s property to yield distinct activation patterns for different architectures. Concretely, every image yields a binary activation vector upon passing through a network, forming a binary matrix for a mini-batch. The logarithm of the determinant of this matrix serves as a scoring metric. Authors show that larger naswot values are associated with better training performances, which leads to the conclusion that high-performing networks should be able to distinguish the inputs before training. Unfortunately, the method can only be implemented on networks with ReLU activation functions, which limits its applicability to convolutional architectures.
@@
Another work employing the aforesaid ReLU property is the one from Chen et al. (2021). They combine the number of linear regions in the input space with the spectrum of the neural tangent kernel (NTK) to build the tenas metric. Instead of evaluating each network in the search space individually, they create a super-network built with all the available edges and operators and then prune it.
@@
Coefficient of variance. Our previous work on fully trainless NAS (Gracheva (2021)) evaluates the stability of untrained scores over random weights initialisations. The networks are initialised with multiple random seeds, and architectures are selected based on the coefficient of variance of the accuracy at initialisation, cv. While cv performance is associated with a high error rate, the author concludes that a good architecture should be stable against random weight fluctuations. While this method can, in theory, apply to any neural architecture type, it requires multiple initialisations and is relatively heavy compared to naswot and later methods. Furthermore, accuracy-based scoring metrics can only apply to classification problems, and it is unclear how to extend cv implementation to the regression tasks.
@@
Gradient sign. The grad_sign metric is built to approximate the sample-wise optimisation landscape (Zhang and Jia (2021)). The authors argue that the closer local minima for various samples sit to each other, the higher the probability that the corresponding gradients will have the same sign. The number of samples yielding the same gradient sign approximates this probability. It allows to evaluate the smoothness of the optimisation landscape and architecture trainability. The method requires labels and gradient computation.
@@
Pruning-at-initialisation proxies. Several powerful zero-cost proxies have emerged as an adaptation of pruning-at-initialisation methods to NAS in the work by Abdelfattah et al. (2021): grad_norm (Wang et al. (2020)), snip (Lee et al. (2018)), synflow (Tanaka et al. (2020)). These metrics are originally developed to evaluate the network’s parameters’ salience and prune away potentially meaningless synapses. They require a single forward-backwards pass to compute the loss. Then, the importance of parameters is computed as a multiplication of the weight value and gradient value. Then, they integrate the salience over all the parameters in the network to evaluate its potential upon training. What is particularly interesting about the synflow metric is that it evaluates the architectures without looking at the data by computing the loss as the product of all the weights’ values (randomly initialised). synflow metric shows the most consistent performance among various search spaces and sets the state-of-the-art for the zero-cost NAS.
@@
Both naswot and synflow do not depend on labels, which arguably reduces the effect of human error during data labelling. Moreover, naswot does not require gradient computation, which renders this method less memory-intensive.
@@
The above results imply that neural networks might have some intrinsic property which defines their prediction potential before training. Such property should not depend on the values of trainable parameters (weights) but only on the network’s topology. In the present work, we combine the takeaways from the existing trainless NAS implementations to present a new metric which significantly outperforms existing zero-cost NAS methods.
1.2 NAS benchmarks
To guarantee our metric’s reproducibility and compare its performance against other NAS algorithms, we evaluate it on the three widely used NAS benchmark datasets.
NAS-Bench-101. The first and one of the largest NAS benchmark sets of trained architectures. It consists of $423{,}624$ cell-based convolutional neural networks (Ying et al. (2019)). The architectures consist of three stacks of cells, each followed by max-pooling layers. Cells may have up to $7$ vertices and $9$ edges, with $3$ possible operations. This benchmark is trained multiple times on a single dataset, CIFAR-10 (Krizhevsky et al. (2009)), with a fixed set of hyperparameters for $108$ epochs.
NAS-Bench-201. It is a set of architectures with a fixed skeleton consisting of a convolution layer and three stacks of cells connected by a residual block (Dong and Yang (2020)). Each cell is a densely connected directed acyclic graph with $4$ nodes, $5$ possible operations and no limits on the number of edges, providing a total of $15{,}625$ possible architectures. Architectures are trained on three major datasets: CIFAR-10, CIFAR-100 (Krizhevsky et al. (2009)) and a downsampled version of ImageNet (Chrabaszcz et al. (2017)). Training hyperparameters are fixed, and the training spans $200$ epochs.
NAS-Bench-NLP. As the name suggests, this benchmark consists of architectures suitable for neural language processing (Klyuchnikov et al. (2022)). Concretely, it consists of randomly-generated recurrent neural networks. Recurrent cells comprise $24$ nodes, $3$ hidden states and $3$ input vectors at most, with $7$ allowed operations. Here, we only consider models trained and evaluated on Penn Treebank (PTB, Marcinkiewicz (1994)) dataset: $14{,}322$ random networks with a single seed and $4{,}114$ with tree seeds. The training spans $50$ epochs and is conducted with fixed hyperparameters.
2 Epsilon metric
@@
Two existing NAS methods inspire the metric that we share in the present work: cv (Gracheva (2021)) and weight agnostic neural networks (Gaier and Ha (2019)). Both metrics aim to exclude individual weight values from consideration when evaluating networks: the former cancels out the individual weights via multiple random initialisations, while the latter sets them to the same value across the network. It is very intriguing to see that an network can be characterised purely by its topology.
@@
As mentioned above, the cv metric has two principal disadvantages. While it shows a fairly consistent trend with trained accuracy, it suffers from high uncertainty. It can be, to some degree, explained by the fact that random weight initialisations bring in some noise. Our idea is that replacing random initialisations with single shared weight initialisations, similarly to Gaier and Ha (2019), should improve the method’s performance.
@@
The second weak point is that cv is developed for classification problems and relies on accuracy. Therefore, it needs to be clarified how to apply this metric to regression problems. The coefficient of variation is a ratio of standard deviation to mean, and Gracheva (2021) shows that cv correlates negatively with train accuracy. It implies that the mean untrained accuracy should be maximised. On the other hand, for regression tasks, performance is typically computed as some error, which is sought to be minimised. It is not apparent whether the division by mean untrained error would result in the same trend for cv metric.
@@
To address this issue, we decided to consider raw outputs. This modification renders the method applicable to any neural architecture. However, it comes with a difference: accuracy returns a single value per batch of data, while raw outputs are $[N_{\textrm{BS}}\times L]$ matrices, where $N_{\rm{BS}}$ is the batch size and $L$ is the length of a single output. [^1] This length depends on the task and architecture: for regression tasks $L=1$, for classification, $L$ is equal to the number of classes, and for recurrent networks, it depends on the desired length of generated string. In our work, we flatten these matrices to obtain a single vector $\mathbf{v}$ of length $L_{v}=N_{\textrm{BS}}\times L$ per initialisation. We then stuck both initialisations into a single output matrix $\mathbf{V}$.
@@
Before proceeding to statistics computation over initialisations, we also must normalise the output vectors: in the case of constant shared weights, outputs scale with weight values. In order to compare initialisations on par with each other, we use min-max normalisation:
@@
$\delta=\frac{1}{L_{v}}\sum\limits_{j=0}^{L_{v}}\lvert\mathbf{V}_{1,j}^{\prime}-\mathbf{V}_{2,j}^{\prime}\rvert.$ (2)
@@
$\mu=\frac{1}{L_{v}}\sum\limits_{j=0}^{L_{v}}\frac{\mathbf{V}_{1,j}^{\prime}+\mathbf{V}_{2,j}^{\prime}}{2}=\frac{1}{2L_{v}}\sum\limits_{i=0}^{2}\sum\limits_{j=0}^{L_{v}}\mathbf{V}_{i,j}^{\prime}$ (3)
Finally, the metric is computed as the ratio of $\delta$ and $\mu$:
$\varepsilon=\frac{\delta}{\mu}.$ (4)
We refer to our metric as epsilon, as a tribute to the $\varepsilon$ symbol used in mathematics to denote error bounds. Algorithm 1 details the epsilon metric computation.
Algorithm 1 Algorithm for epsilon metric computation
@@
     Compute epsilon metric (Equation 4)
@@
2.1 Empirical evaluation
Here we evaluate the performance of epsilon and compare it to the results for zero-cost NAS metrics reported in Abdelfattah et al. (2021). We use the following evaluation scores (computed with NaN omitted):

### v3

# Neural Architecture Search: Two Constant Shared Weights Initialisations

@@ Abstract
In the last decade, zero-cost metrics have gained prominence in neural architecture search (NAS) due to their ability to evaluate architectures without training. These metrics are significantly faster and less computationally expensive than traditional NAS methods and provide insights into neural architectures’ internal workings. This paper introduces epsinas, a novel zero-cost NAS metric that assesses architecture potential using two constant shared weight initialisations and the statistics of their outputs. We show that the dispersion of raw outputs, normalised by their average magnitude, strongly correlates with trained accuracy. This effect holds across image classification and language tasks on NAS-Bench-101, NAS-Bench-201, and NAS-Bench-NLP. Our method requires no data labels, operates on a single minibatch, and eliminates the need for gradient computation, making it independent of training hyperparameters, loss metrics, and human annotations. It evaluates a network in a fraction of a GPU second and integrates seamlessly into existing NAS frameworks. The code supporting this study can be found on GitHub at https://github.com/egracheva/epsinas.
keywords:
machine learning, neural architecture search, zero-cost NAS, efficient NAS, NAS benchmarks
@@ 1 Introduction
The field of neural architecture search (NAS) emerged roughly a decade ago as a solution to automate the process of neural geometry optimisation. In its early stages, NAS predominantly relied on evaluating candidate architectures through computationally intensive training processes, using techniques such as reinforcement learning [49], evolutionary algorithms [38, 31], and Bayesian optimisation [13, 47].
One-shot algorithms speed up the search by sharing weights instead of training each architecture separately. These methods include efficient reinforcement learning [37], random search with parameter sharing [27], and differentiable approaches [30, 7, 50]. However, they require training a large hypernetwork, necessitating intricate hyperparameter tuning. While these techniques are efficient, they often fail to deliver consistent results [11]. For example, even advanced approaches like DARTS- [7] exhibit significant variability compared to evolutionary or reinforcement learning methods.
Some methods attempt to estimate network performance without training on the dataset of interest by leveraging an auxiliary predictive machine learning (ML) model built on a dataset of pre-evaluated architectures [19, 9]. Although these methods accelerate the NAS process for image recognition, they still rely on training and are not applicable to other domains of ML.
@@
2 Related work
To simplify architecture search, zero-cost NAS methods identify optimal architectures without requiring full training. These methods leverage zero-cost proxies [16, 34, 14, 18, 25], typically demanding computational resources equivalent to just one or a few training epochs. Consequently, zero-cost NAS is two to three orders of magnitude faster than traditional NAS approaches. Below, we summarize the most prominent zero-shot NAS methods. For a comprehensive field review, see [26].
[…]
@@ 2 Related work
ReLU activation pattern. In 2020, [34] introduced the naswot metric, which exploits the property of the rectified linear unit (ReLU) activation function [2] to produce distinct activation patterns for different architectures. Specifically, every image passing through a network generates a binary activation vector, which forms a binary matrix for a mini-batch. The logarithm of the determinant of this matrix serves as a scoring metric. The authors demonstrated that higher naswot values are associated with better training performance, concluding that high-performing networks should be able to distinguish inputs even before training. However, this method is limited to networks using ReLU activation functions, which restricts its applicability to convolutional architectures.
@@
A conceptually similar work by [4] combines the number of linear regions in the input space with the spectrum of the neural tangent kernel (NTK) to develop the tenas metric. Instead of individually evaluating each network in the search space, they construct a super-network that includes all the available edges and operators, which is then pruned.
@@
Coefficient of variance. Our previous work on fully trainless NAS [18] evaluated the stability of untrained scores across random weight initialisations. Networks were initialised with multiple random seeds, and architectures were selected based on the coefficient of variance of the accuracy at initialisation, cv. While cv performance is associated with a high error rate, the study concluded that a good architecture should demonstrate stability against random weight fluctuations. Although this method theoretically applies to any type of neural architecture, it requires multiple initialisations and is relatively computationally expensive compared to naswot and later methods. Additionally, accuracy-based scoring metrics are limited to classification problems, and it remains unclear how to adapt the cv approach for regression tasks.
Inverse variance of gradients. A recent study by [25] introduced the zico metric, a zero-cost proxy for neural architecture performance that, like cv [18], is derived from network statistics. zico is computed as the ratio of the average gradient values to their standard deviation—in essence, a reversed cv, where accuracy is substituted with gradient values. Notably, zico exhibits a strong correlation with trained accuracy, surpassing nparams—the total number of parameters in a network—as a performance estimator. While nparams may seem like a simplistic proxy, it remained the most effective predictor of network performance until 2023.
Gradient signal-to-noise ratio. In 2023, [44] introduced the gradient signal-to-noise ratio metric ($\xi$-gsnr), which, as the name suggests, closely resembles zico. Their approach builds on zico by squaring the metric and incorporating a small constant $\xi$ in the denominator. This seemingly minor modification significantly impacts the metric’s performance, establishing a new state-of-the-art zero-cost NAS. Additionally, the authors provide essential theoretical justification for their approach. A comparison of the cv, zico, and $\xi$-gsnr methodologies is presented in Table 1.
@@
Gradient sign. The grad_sign metric is designed to approximate the sample-wise optimisation landscape [52]. The authors argue that when local minima for various samples are closer, the probability increases that the corresponding gradients will have the same sign. The number of samples yielding the same gradient sign approximates this probability, allowing for the smoothness evaluation of the optimisation landscape and the architecture’s trainability. However, this method requires labels and gradient computation.
@@
Pruning-at-initialisation proxies. Several promising zero-cost proxies have emerged as adaptations of pruning-at-initialisation methods for NAS, as demonstrated in the work by [1]: grad_norm [46], snip [24], and synflow [45]. These metrics were originally developed to assess the salience of a network’s parameters and prune potentially irrelevant synapses. They require a single forward-backwards pass to compute the loss, after which the importance of parameters is calculated as the product of the weight and gradient values. The salience is then integrated over all parameters in the network to estimate its potential performance once trained.
What makes the synflow metric particularly notable is its ability to evaluate architectures without referencing the data, as it computes the loss based solely on the product of all randomly initialised weights’ values. Among these proxies, the synflow metric demonstrates the most consistent performance across various search spaces and sets the state-of-the-art for zero-cost NAS.
@@
Neither naswot nor synflow depend on labels, which reduces the effect of human error during data labelling. Moreover, naswot does not require gradient computation, which renders this method less memory-intensive.
A few other efficient zero-shot methods cannot be directly compared to our results (such as Zen-NAS by [28]). For more details on the current status of NAS, please refer to the most recent systematic review on NAS [40].
@@
The results of the above studies suggest that neural networks possess an intrinsic property that determines their predictive potential even before training. This property should be independent of the specific values of trainable parameters (weights) and instead depend solely on the network’s topology. In this work, we build upon insights from existing zero-cost NAS implementations to propose a new metric that shows strong performance and outperforms current zero-cost NAS methods.
3 Epsinas metric
@@
Two existing NAS methods inspire the metric that we share in the present work: cv [18] and weight agnostic neural networks [16]. Both metrics aim to reduce the role of individual weights in the network to access its topology. The cv metric levels out the individual weights via multiple random initialisations, while [16] sets the weights to the same value across the network.
@@
As mentioned, the cv metric has two principal disadvantages. While it shows a fairly consistent trend with trained accuracy, it suffers from high statistical deviations. To some degree, it must be due to the noise from random weight initialisations. With epsinas, we replaced random initialisations with single shared weight initialisations to improve the method’s performance.
@@
The second weak point of cv is that it has been developed for classification problems and relies on accuracy, which should be maximised. On the other hand, for regression tasks, performance is typically computed as an error, which is sought to be minimised. The cv is a ratio of standard deviation and the mean of untrained accuracies, and the final metric correlates negatively with train accuracy. It is not apparent whether the division by mean untrained error would result in the same trend for cv metric. To address this issue, we decided to step out from evaluation metrics and consider raw outputs instead. This modification renders the method applicable to any neural architecture.
@@
For this, we flatten the raw output matrices to obtain a single vector $\mathbf{v}$ of length $L_{\mathbf{v}}=N_{\textrm{BS}}\times L_{\mathbf{O}}$ per initialisation, where $N_{\rm{BS}}$ is the batch size and $L_{\mathbf{O}}$ is the length of a single output  [^1] This length depends on the task and architecture: for regression tasks $L_{\mathbf{O}}=1$, for classification, $L_{\mathbf{O}}$ is equal to the number of classes, and for recurrent networks, it depends on the desired length of generated string..
@@
Before proceeding to statistics computation over initialisations, we also must normalise the output vectors: in the case of constant shared weights, outputs scale with weight values. To compare initialisations on par with each other, we use min-max normalisation:
@@
$\mathrm{MAE}=\frac{1}{L_{\mathbf{v}}}\sum\limits_{j=0}^{L_{\mathbf{v}}}\lvert\mathbf{v}_{0,j}^{\prime}-\mathbf{v}_{1,j}^{\prime}\rvert.$ (2)
@@
$\mu=\frac{1}{L_{\mathbf{v}}}\sum\limits_{j=0}^{L_{\mathbf{v}}}\frac{\mathbf{v}_{1,j}^{\prime}+\mathbf{v}_{2,j}^{\prime}}{2}$ (3)
Finally, the metric is computed as the ratio of $\mathrm{MAE}$ and $\mu$:
$\varepsilon=\frac{\mathrm{MAE}}{\mu}.$ (4)
We use real data from the respective datasets to compute the statistics, as suggested in [14]. In Section 6.5, we demonstrate that using real data enhances search efficiency.
We refer to our metric as epsinas, paying tribute to the $\varepsilon$ symbol commonly used in mathematics to denote error bounds. The epsinas metric was developed independently but in parallel with zico and follows a similar approach. While zico uses the mean and standard deviation of gradients over two mini-batches, our metric combines the mean and mean absolute error of raw network outputs. This parallel development highlights a shared idea — leveraging statistical properties of untrained networks to guide architecture selection.
Table 1 compares epsinas to the three the most resembling zero-cost proxies: cv, zico, and $\xi$-gsnr. Although the difference between the two methods may seem minor and zico and $\xi$-gsnr show a strong correlation with trained accuracy, epsinas has an advantage: it does not require backpropagation for gradients computation, making it faster to calculate. Additionally, as we demonstrate in the next section, epsinas exhibits an even stronger correlation with trained accuracy.
Table 1: Comparison of epsinas to the three the most resembling zero-cost proxies. Data $\mathbf{D}$ denotes the data used for metrics computation. $N_{init}$ refers to the number of initialisations (for zico and $\xi$-gsnr we report the number of batches).
Metric Data, $\mathbf{D}$ Formula Initialisation $N_{init}$ Gradients True labels
cv Accuracy $\frac{\sigma(\mathbf{D})}{\mu(\mathbf{D})}$ Kaiming normal 100 No Yes
zico Gradients $\frac{\mu(\mathbf{D}}{\sigma(\mathbf{D})}$ Kaiming normal 2 Yes Yes
$\xi$-gsnr Gradients $\frac{\mu(\mathbf{D})^{2}}{\sigma(\mathbf{D})^{2}+\xi}$ Kaiming normal 8 Yes Yes
epsinas Normalised outputs $\frac{\mathrm{MAE}(\mathbf{D})}{\mu(\mathbf{D})}$ Constant shared weights 2 No No
Algorithm 1 details the epsinas metric computation.
Algorithm 1 Algorithm for epsinas metric computation
@@
     Compute epsinas metric (Equation 4)
@@
Proper initialization is crucial for implementing the metric effectively. In our current implementation, we preserve the initialization of biases, batch normalization, and embeddings. Specifically, bias weights are set to zero, batch normalization parameters are initialized to ones, and embeddings follow the original paper’s approach using random integers [20].
4 Experimental design
4.1 NAS benchmarks
Following standard NAS metric evaluation practices, we assess epsinas on three well-established NAS benchmark datasets containing fully trained neural architectures. To ensure the broad applicability of epsinas across different architectures and tasks, we evaluate it on both image classification and language processing tasks.
NAS-Bench-101. The first and one of the most extensive NAS benchmarks, NAS-Bench-101, consists of $423{,}624$ fully trained convolutional neural networks [51]. Each architecture comprises three stacked cells, followed by max-pooling layers, with up to $7$ vertices, $9$ edges, and $3$ possible operations per cell. All models are trained on CIFAR-10 [21] for $108$ epochs using fixed hyperparameters.
NAS-Bench-201. This benchmark provides $15{,}625$ architectures with a fixed backbone: a convolutional layer followed by three stacked cells connected by a residual block [12]. Each cell is a densely connected directed acyclic graph with $4$ nodes and $5$ possible operations. Architectures are trained on CIFAR-10, CIFAR-100 [21], and a downsampled version of ImageNet [5], all with fixed hyperparameters over $200$ epochs.
NAS-Bench-NLP. Recurrent cells comprise $24$ nodes, $3$ hidden states and $3$ input vectors at most, with $7$ allowed operations. Here, we only consider models trained and evaluated on Penn Treebank (PTB) data [33]: $14{,}322$ random networks with a single seed. The training spans $50$ epochs and is conducted with fixed hyperparameters.
4.2 Evaluation metrics
We evaluate the performance of epsinas and compare it to the results for zero-cost NAS metrics reported in [1]. We use the following evaluation scores (computed with NaN omitted):